Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
menzione - 16:04 Wednesday 22 June 2022 (56224) Print this report
Operator Report - Morning shift

Upon arrival, I found ITF in relocking sequence. I lock it on IR then the planned ISC activity on Autorelock Metatron node (Bersanetti, Carbognani) started and went on without major problems for the whole shift.

Parallel activities:
- ENV - 12:00 to 12:30 UTC - under my suggestion, Derosa went to CB in order to replace the Temperature probe. 
- CAL - 11:00 UTC - Photon Calibrator checks in CB, WE, and NE (Lapp team)

Sub-system reports

Air Conditioning
ENV..EDB_HU out of threshold. (DET_Area red flag on MS).
Experts informed and the task force is looking on it. 13:40 UTC the humidity is now around 50% and on EDB the noise is higher than before, ACC_Z noisy (plot1)

At 11:45 UTC the spot of B1p was in the lower part of the screen. With Diego we fixed the position via VPM (SDB1_LC):
- Used fixed setpoints (spot again in the center of the screen)
- Open B5 beam drift
- Close B5 beam drift
- Use floating setpoints
- Open B5 beam drift

06:45 UTC - restored the vertical position and closed properly SQB1 loop.
10:00 UTC - FCIM_LC drift of all DOF. M. DeLaurenti asked Bertolini to restore the standard state.

09:30, 12:33 UTC - WE loop opened by the Guardian. Loop properly closed and realigned.

Images attached to this report
Search Help